OpenStudy (anonymous):

simplify 2 2 -- x -- 3a a^2

OpenStudy (lalaly):

did u try to do it?

OpenStudy (anonymous):

Do I cross multiply? (I dont always pay attention in math......)

OpenStudy (lalaly):

when multiplying fractions u multiply top by top and bottom by bottom

OpenStudy (lalaly):

u cross multiply when there is an equal sign between them,, but when there is a multiplication sign u simply multiply top by top and then bottom by bottom

OpenStudy (anonymous):

But what if its something like in the problem when the denominator is a variable to the power of 2? does the variable become one?

OpenStudy (lalaly):

when both variabls are the same,, like in this question u have 3a and a^2 that becomes 3a^(2+1)=3a^3 but if they were different like for example a^2 and b u just write it as a^2b

OpenStudy (anonymous):

oh, ok. had to think about that for a second.

OpenStudy (lalaly):

so now do u know how to find the product of these fractions?

OpenStudy (anonymous):

I think so. would the answer be 4 over 3a^3?

OpenStudy (lalaly):

yup

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

\[2/3a x2/3a ^2=4/3a ^3\]

OpenStudy (anonymous):

Thanks pastory, but I had just figured it out before you posted it. :D

OpenStudy (anonymous):

thanks